# Magnetic Field-Line Curvature and Its Role in Particle Acceleration by Magnetically Dominated Turbulence

### v1

@@ 1 Introduction
In this work, we focus on how particles gain energy from the motional electric field induced by turbulent plasma motions, with particular attention to acceleration via curvature-drift motion along the motional electric field. PIC simulations suggest that shear of the motional electric field along magnetic field lines is important for reconstructing particle energy histories (Bresci et al., 2022), yet the curvature-drift contribution in turbulence remains largely unexplored. Here, we employ a guiding-center framework to provide a quantitative assessment of this contribution and its dependence on $\delta B_{0}/B_{0}$. While guiding-center studies have extensively explored curvature-drift acceleration in magnetic reconnection (Drake et al., 2006; Dahlin et al., 2014; Guo et al., 2014, 2015; Li et al., 2015, 2017; Dahlin et al., 2017; Rowan et al., 2019), where ideal electric fields are shaped by large-scale reconnection outflows, its role in turbulent plasmas is far less clear. In turbulence, motional electric fields are intermittent and disordered, and the tangled magnetic geometry creates numerous potential sites for curvature-drift acceleration. Understanding this process is key to explaining how turbulence transfers energy to nonthermal particles and to linking kinetic-scale plasma physics with observable astrophysical signatures.
@@ 3 Turbulence Statistics
The statistical properties of magnetic field-line curvature depend strongly on the fluctuation-to-mean magnetic field ratio, as illustrated in Fig. 5. Panel (a) shows the PDFs of $\kappa l_{c}$ for simulations with different $\delta B_{0}/B_{0}$. The distributions are broad, with power-law tails at both low and high curvatures. For reference, power-law slopes of $\kappa^{+1}$ and $\kappa^{-5/2}$ are indicated as asymptotic scalings for low- and high-curvature limits, respectively (see next paragraph). As the mean magnetic field decreases, the peaks of the PDFs shift to higher curvatures, as $\kappa_{\rm rms}\sim l_{c}^{-1}\delta B_{0}/\sqrt{B_{0}^{2}+\delta B_{0}^{2}}$, and the distributions extend to $\kappa$ values much larger than the peak, reflecting the strong intermittency of the magnetic field, with localized regions of intense curvature forming within more tangled magnetic configurations. [^3] We verified that the curvature PDFs remain unchanged when measured in either the simulation frame or the ${\bm{E}}\times{\bm{B}}$ frame.
@@
where $\lambda>0$ is the scale parameter and $\alpha>0$ is the shape parameter. This model reproduces the probability density functions well, with fitted values $\lambda=2.5,7,9$ for $\delta B_{0}/B_{0}=1/3,1,3$, respectively, reflecting the increasing “typical” strength of stochastic field-line contraction events as the mean field weakens. The corresponding shape parameters are $\alpha=8,5,3$, indicating that extreme field-line contraction events become more frequent under weaker mean-field conditions. Thus, large-amplitude turbulence not only raises the characteristic amplitude of field-line contraction (through larger $\lambda$) but also enhances intermittency (through smaller $\alpha$), making stochastic energization more efficient.
@@ Acknowledgments
We thank James Beattie, Philipp Kempski, and Emanuele Sobacchi for fruitful discussions. This work was supported by the NSF Research Experiences for Undergraduates (REU) program under grant NSF PHY-2447137. L.C. acknowledges support from NSF PHY-2308944 and NASA ATP 80NSSC24K1230. Additional support was provided in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR) contract C090171.

### v2

@@ 1 Introduction
In this work, we focus on how particles gain energy from the motional electric field induced by turbulent plasma motions, with particular attention to acceleration via curvature-drift motion along the motional electric field. PIC simulations suggest that shear of the motional electric field along magnetic field lines is important for reconstructing particle energy histories (Bresci et al., 2022), and analytical frameworks have discussed in detail the different forces acting on the particles (Lemoine, 2021, 2022, 2025). Here, we employ a guiding-center framework to provide a quantitative assessment of curvature-drift acceleration and its dependence on $\delta B_{0}/B_{0}$. Guiding-center studies have extensively explored curvature-drift acceleration in magnetic reconnection (Drake et al., 2006; Dahlin et al., 2014; Guo et al., 2014, 2015; Li et al., 2015, 2017; Dahlin et al., 2017; Rowan et al., 2019), where ideal electric fields are shaped by large-scale reconnection outflows. In turbulence, motional electric fields are intermittent and disordered, and the tangled magnetic geometry creates numerous potential sites for curvature-drift acceleration. Understanding this process is key to explaining how turbulence transfers energy to nonthermal particles and to linking kinetic-scale plasma physics with observable astrophysical signatures.
@@ 3 Turbulence Statistics
The statistical properties of magnetic field-line curvature depend strongly on the fluctuation-to-mean magnetic field ratio, as illustrated in Fig. 5. Panel (a) shows the PDFs of $\kappa l_{c}$ for simulations with different $\delta B_{0}/B_{0}$. The distributions are broad, with power-law tails at both low and high curvatures. For reference, power-law slopes of $\kappa^{+1}$ and $\kappa^{-5/2}$ are indicated as asymptotic scalings for low- and high-curvature limits, respectively (see next paragraph). As the mean magnetic field decreases, the peaks of the PDFs shift to higher curvatures, as $\kappa_{\rm rms}\sim l_{c}^{-1}(1+B_{0}^{2}/\delta B_{0}^{2})^{-1}$, and the distributions extend to $\kappa$ values much larger than the peak, reflecting the strong intermittency of the magnetic field, with localized regions of intense curvature forming within more tangled magnetic configurations. [^3] We verified that the curvature PDFs remain unchanged when measured in either the simulation frame or the ${\bm{E}}\times{\bm{B}}$ frame.
@@
where $\lambda>0$ is the scale parameter and $\alpha>0$ is the shape parameter. This model reproduces the probability density functions well, with fitted values $\lambda=2.5,7,9$ for $\delta B_{0}/B_{0}=1/3,1,3$, respectively, reflecting the increasing “typical” strength of stochastic field-line contraction events as the mean field weakens. The corresponding shape parameters are $\alpha=8,5,3$, indicating that extreme field-line contraction events become more frequent under weaker mean-field conditions. Thus, large-amplitude turbulence not only raises the characteristic amplitude of field-line contraction (through larger $\lambda$) but also enhances intermittency (through smaller $\alpha$), making stochastic energization more efficient ($\dot{\gamma}/\gamma\propto{\bm{v}}_{E}\cdot{\bm{\kappa}}$).
@@ Acknowledgments
We thank James Beattie, Philipp Kempski, and Emanuele Sobacchi for fruitful discussions. We are grateful to Yue Hu for providing the script used to compute the structure functions. This work was supported by the NSF Research Experiences for Undergraduates (REU) program under grant NSF PHY-2447137. L.C. acknowledges support from NSF PHY-2308944 and NASA ATP 80NSSC24K1230. Additional support was provided in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR) contract C090171.